OpenStudy (anonymous):

how to do a substitution on 5x-y=-13,-7x-3y=5

OpenStudy (anonymous):

i know how to do this um i'm tired i'm sorry i think it's best for you to subtract the y's

OpenStudy (anonymous):

i tried but the answer wont come out right

OpenStudy (nikato):

@holla! u can try to isolate the y in the first equation

OpenStudy (anonymous):

http://www.youtube.com/watch?v=-fZLAYn9sGY this is the obvious answer 2 diss

OpenStudy (anonymous):

well when I pluged it in it came to (-2,3) which is right I juz don't know how to solve it to get that answer

OpenStudy (phi):

start with the first equation 5x-y=-13 add + y to both sides, like this 5x -y + y = -13 + y on the left side, what does -y + y simplify to ?

OpenStudy (anonymous):

so what are you cheating on?

OpenStudy (anonymous):

well iv been trying to do this equation for quit awhile and I cant seem to get the answer to come out to (-2,3)

OpenStudy (anonymous):

let me guess you just made this today

OpenStudy (anonymous):

ur account

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

help /_\

OpenStudy (phi):

add + y to both sides, like this 5x -y + y = -13 + y on the left side, what does -y + y simplify to ?

OpenStudy (phi):

what is 1 y take away 1 y ?

OpenStudy (anonymous):

o so okY THIS IS WHAT I PUT NOW 5x=-13+y

OpenStudy (phi):

now add +13 to both sides of 5x=-13+y 5x+13= -13+13+y simplify the right side

OpenStudy (phi):

you should get 5x+13 = y or (the same thing) y= 5x+13 think of this as y is "short-hand" for (5x+13) use the "long version" of y in the second equation -7x-3y=5 that means erase the y, and put (5x+13) in its place. can you do that ?

OpenStudy (anonymous):

I got it now! got the answer thanks!:)